OpenStudy (anonymous):

1. A new cruise ship line has just launched 3 new ships. The Pacific Paradise, the Caribbean Paradise, & the Mediterranean Paradise. The Caribbean Paradise has 37 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 40 fewer deluxe staterooms than 4 times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ship if the total number of deluxe staterooms for the 3 ships is 819.

OpenStudy (anonymous):

Let x= number of deluxe staterooms in Pacific Paradise, y= number of deluxe staterooms in Caibbean Paradise, and z= number of deluxe staterooms in Mediterranean Paradise. As you read each sentence, your equations should look like this: y= 37+x z=4*x-40 x+y+z=819 So we now need to substitute: x+(37+x)+(4x-40)=819 6x-3=819 6x=822 x=137 So there are 137 rooms on the Pacific Paradise, 174 rooms on the Caribbean Paradise (137+37), and 508 rooms on Mediterranean Paradise(4*137-40).